OpenStudy (anonymous):

f(x) = 8 – 105 x f(35) =

OpenStudy (onepieceftw):

This is gibberish...

OpenStudy (onepieceftw):

F(x)=8-105x?

OpenStudy (anonymous):

105 over x

OpenStudy (anonymous):

f(x) = 8 - 105 over x f (35)=

OpenStudy (onepieceftw):

\[F(x)=8-\frac{ 105 }{ x }\] Alright, so substitute x for 35

OpenStudy (anonymous):

5 i think

OpenStudy (anonymous):

@OnePieceFTW

OpenStudy (onepieceftw):

Yes, that's it.

OpenStudy (anonymous):

Ight thank you @OnePieceFTW

OpenStudy (anonymous):

hunter @i<3u43v4

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

who

OpenStudy (anonymous):

no i didnt idk who u talking about

OpenStudy (anonymous):

its not even june yet and i den have a account last year

OpenStudy (anonymous):

@stoner98